# Two dimensional dimers beyond planarity

### v2

@@ 5 Probabilistic estimates
For the first identity below we use Proposition 3.1, for the third inequality we use Proposition 4.5, obtaining that
$\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>M\big{)}$ $\leq\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}-m_{\{o,g\}}^{2}>M\big{)}$
$\leq\sum\limits_{u=M+1}^{\infty}\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}-m_{\{o,g\}}^{2}=u\big{)}$
$=\sum\limits_{u=M+1}^{\infty}\langle\,g_{u,o}(s_{o})\,\big{\rangle}$
$\leq\sum\limits_{u=M+1}^{\infty}\Big{(}\langle\prod_{x\in V}g_{u,x}(s_{x})\big{\rangle}\Big{)}^{\frac{1}{|V|}}$
$=\sum\limits_{u=M+1}^{\infty}\mathcal{P}_{L,K,\beta,h,r}\big{(}\forall x\in V:n_{x}^{i}-m_{\{x,g\}}^{2}=u\big{)}^{\frac{1}{|V|}}$
$\leq\sum\limits_{u=M+1}^{\infty}\Big{(}\frac{1}{Z_{L,K,\beta,h,r}}d^{|V|}\frac{\max\{\beta,1\}^{u|V|}\,e^{2(|E|+|E_{g}|)\max\{\beta,1\}}}{\big{(}(u/d)!\big{)}^{\frac{1}{d}}}\Big{)}^{\frac{1}{|V|}}$
$\leq c\sum\limits_{u=M+1}^{\infty}\frac{1}{\big{(}(u/d)!\big{)}^{\frac{1}{d}}}$
For the last inequality we used a crude upper bound for the probability of the event $\{\forall x\in V:n_{x}-m_{\{x,g\}}^{2}=u\}$ by using the fact that, if $n_{x}(w)-m_{\{x,g\}}^{2}(w)=u$, then there exists one of the $d$ edges of $E$ touching $x$ with at least $u/d$ dimers of colour $1$ and there are at most $d$ possibilities for choosing such edge, and for each remaining edge and each colour we summed the number of dimers independently (using $U\leq 1$ and ignoring the constraint on the dimer cardinality) getting a factor $e^{\beta}$. For the last inequality we also used that $Z_{L,K,\beta,h,r}\geq c^{|V|}$ for some $c=c(\beta)$. Being the sum convergent we can make it arbitrarily small choosing $M$ large, obtaining (5.3).
@@
For (5.4) we observe that if $U(0)=0$ then the proof is obvious by definition of the measure since by our assumption there exists $k>0$ such that $U(k)>0$. Hence, we can assume that $U(0)>0$. By Proposition 3.1 and Proposition 4.5 we obtain that
$\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}=0\big{)}$ $=\langle g_{o}(s_{o})\big{\rangle}$
$\leq\Big{(}\langle\prod_{x\in V}g_{x}(s_{x})\big{\rangle}\Big{)}^{\frac{1}{|V}|}$
$=\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}\forall x\in V,n_{x}=0\big{)}\Big{)}^{\frac{1}{|V|}}$
$=\Big{(}\frac{U(0)^{|V|}}{Z_{G,\beta,h,r}}\Big{)}^{\frac{1}{|V|}}<c<1.$
To show that $c<1$ we identify a subset $A\subset V$ such that $|A|>\frac{1}{100}|V|$ and $d(x,y)>2$ for each distinct $x,y\in A$. We then lower bound $Z_{G,\beta,h,r}$ by considering only the contribution of the configurations which have the vertices of $A$ either empty or touched by precisely $k$ blue and $k$ red dimers on the same edge and all the remaining vertices empty, hence obtaining $Z_{G,\beta,h,r}\geq U(0)^{|V|}(1+\frac{\beta^{2k}U(k)^{2}}{k!^{2}U(0)^{2}})^{{\frac{|V|}{100}}}$. This leads to the desired bound and concludes the proof. ∎
@@
For any $\beta>0$, $h\in[0,1]$, $r\in\mathbb{N}$, there exists a constant $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$, for any $z\sim o$, we have that,
$\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle_{L,K,\beta,h,r}\,\,\,$ $=\frac{1}{\beta}\mathcal{E}_{\mathbb{S}_{L,K},\beta,h,r}\big{(}\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big{)}\leq\frac{1}{\beta},$ (5.5)
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle$ $\geq c.$ (5.6)
@@
$\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle_{L,K,\beta,h,r}=\frac{\mu_{\beta,h,r}(\mathcal{A})}{\mathcal{Z}_{G,\beta,h,r}},$
where $\mathcal{A}$ is the set of configurations with precisely one $i$-dimer on $\{o,s\}$, one $i$-dimer on $\{z,s\}$, and no further dimer touching $s$.
We introduce a map $f$ which takes any configuration $(m,\pi)\in\mathcal{A}$ and associates to it the configuration $f(m,\pi)\in\mathcal{W}^{\ell}$ which is defined by following the next three steps:
@@
    adding a $i$-dimer on $\{o,z\}$ and assigning to it the label $m^{i}_{\{o,z\}}+1$,
@@
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle=\sum\limits_{z\sim o}\frac{1}{\beta}\mathcal{E}_{L,K,\beta,h,r}\big{(}\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big{)}\geq\sum\limits_{z\sim o}\frac{1}{M\beta}\mathcal{P}_{L,K,\beta,h,r}\big{(}0<m_{\{o,z\}}^{i},n_{o}^{i}\leq M\big{)}\\ \geq\sum\limits_{z\sim o}\frac{1}{M\beta}\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>0\big{)}-\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>M\big{)}\Big{)}\\ \geq\frac{1}{M\beta}\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>0\big{)}-6\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>M\big{)}\Big{)}\geq c.$
where the last inequality follows from Lemma 5.2 by choosing $M$ large enough depending on $\beta$.
We now introduce an important quantity of interest, the magnetisation,
@@
Set $\beta\geq 0$, $r\in\{1,2\}$, $\ell\in\mathbb{N}$. There exist a constant $c=c(\beta,\ell)\in(0,\infty)$ and a value $h_{0}>0$ small enough such that for any sufficiently small $\tilde{h}\in(0,h_{0})$, every $L,K\in 2\mathbb{N}\cup\{1\}$
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)$ $\geq\,\,\frac{\beta\tilde{h}}{|\mathbb{S}_{L,K}|}\sum\limits_{x\in\mathbb{S}_{L,K}}\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle_{L,K,\beta,0,0},$ (5.8)
$\langle\cos(\ell s_{o}^{1})\rangle_{L,K,\beta,h_{L,K},r}$ $\leq c\,\tilde{h}$ (5.9)
@@
From a Taylor expansion of the exponentials in (2.4) and Lemma 3.2 we deduce that
@@
For a lower bound we use the positive goodness condition (introduced right after Proposition 3.1) and reduce the sum in (5.10) to the terms such that $m_{x}=1$ for just one $x\in V$ and $m_{y}=0$ for all $y\neq x$, thus obtaining (5.8) using the definition of $h_{L,K}$ and that $\ell=r$. For (5.9) we start from (5.10) and use Lemma 5.1 to deduce that
@@
where in the last step we used that $\tilde{h}$ is small enough depending on $\beta$. ∎
6 Cauchy-Schwarz inequality
This section presents the central argument for our new proof of the Merming Wagner theorem for non positive (but reflection positive) measures.
@@ 6 Cauchy-Schwarz inequality
$\tilde{H}:=H+\sum\limits_{x\in\mathbb{S}^{e}_{L,k}}\log(g_{x})+\sum\limits_{x\in\mathbb{S}^{o}_{L,k}}\log(\overline{g_{x}}),$
@@
$A:=(\sin(rs_{x}^{1}))_{x\in\mathbb{S}_{L,K}}\quad\quad B:=\beta\,\big{(}\partial_{s_{x}^{1}}\tilde{H}-\partial_{s_{x}^{2}}\tilde{H}\big{)}_{x\in\mathbb{S}_{L,K}},$
@@
[…]
@@
$\beta^{2}\langle$ $\big{(}\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H}\big{)}\,\big{(}\partial_{s_{y}^{1}}\tilde{H}-\partial_{s_{y}^{2}}\tilde{H}\big{)}\rangle=-\beta\langle\big{(}\partial_{s_{y}^{1}}-\partial_{s_{y}^{2}}\big{)}\,\big{(}\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H}\big{)}\rangle$
=$=\begin{cases}-\beta\,\langle\partial^{2}_{s_{o}^{1}}\tilde{H}+\partial^{2}_{s_{o}^{2}}\tilde{H}\big{)}\rangle&\mbox{ if $y=o$}\\ -\beta\langle\partial_{s_{y}^{1}}\partial_{s_{o}^{1}}\tilde{H}+\partial_{s_{y}^{2}}\partial_{s_{o}^{2}}\tilde{H}\big{)}\rangle&\mbox{ if $y\sim o$}\\ 0&\mbox{ otherwise}\\ \end{cases}$
=$=\begin{cases}\beta\,\sum\limits_{y\sim o}\Big{(}\langle e^{i(s_{o}^{1}-s_{y}^{1})}\rangle+\langle e^{-i(s_{o}^{2}-s_{y}^{2})}\rangle+2hr^{2}\langle\cos(rs_{o}^{1})\rangle\Big{)}&\mbox{ if $y=o$}\\ -\beta\big{(}\langle e^{i(s_{o}^{1}-s_{y}^{1})}\rangle+\langle e^{-i(s_{o}^{2}-s_{y}^{2})}\rangle\big{)}&\mbox{ if $y\sim o$}\\ 0&\mbox{ otherwise. }\\ \end{cases}$
@@ 6.1 Positivity of the Fourier transform
We next state an inequality which is a discrete analogue of the fact that the Fourier transform of a convex function is positive. The bound will be applied to two-point functions (4.12). Since our convexity property (4.13) fails at a neighbourhood of zero, we obtain a bound which may be negative, but which is negligible comparing to the volume $|\mathbb{S}_{L,K}|$. We define for any $M\in\mathbb{N}$, $\mathbb{Z}^{M}_{odd}:=(2\mathbb{Z}+1)\cap(-M/2,M/2]$ and $\mathbb{Z}^{M}_{even}:=(2\mathbb{Z})\cap(-M/2,M/2]$.
@@
    $G(x)=G(-x)$ for any $x\in e_{1}\mathbb{Z}^{L}_{odd}\cup e_{2}\mathbb{Z}^{L}_{odd}\cup e_{3}\mathbb{Z}^{K}_{odd}$,
@@
$\sum_{x\in e_{i}\mathbb{Z}_{odd}^{L}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{i})-G(5e_{i}))\frac{1-\cos 3k_{i}}{2\sin^{2}k_{i}},$ (6.3)
$\sum_{x\in e_{3}\mathbb{Z}_{odd}^{K}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{3})-G(5e_{3}))\frac{1-\cos 3k_{3}}{2\sin^{2}k_{3}},$ (6.4)
@@
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{(2\sin k_{i})^{2}}\sum_{x\in\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x)$ (6.5)
$\sum_{x\in e_{3}\mathbb{Z}^{K}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{(2\sin k_{3})^{2}}\sum_{x\in\mathbb{Z}^{K}_{odd}}(2G(x)-G(x-2e_{3})-G(x+2e_{3}))\cos(k\cdot x).$ (6.6)
@@
$=\frac{1}{(2\sin k_{i})^{2}}\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x),$
@@
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $\geq\frac{1}{(2\sin k_{1})^{2}}\big{(}2(G(5e_{i})-G(3e_{i}))+2(2G(3e_{i})-G(e_{i})-G(5e_{i}))\cos(3k_{i})$
@@
$=\frac{1}{2(\sin k_{1})^{2}}\left((G(e_{i})-G(3e_{i}))(\cos k_{i}-\cos 3k_{i})+(G(3e_{i})-G(5e_{i}))(\cos 3k_{i}-1)\right),$
@@
6.2 Statement and proof of the Cauchy-Schwarz inequality
We now state a Cauchy-Schwarz inequality for the quantities $A$ and $B$ which have been defined above. Our inequality holds for small external magnetic fields, namely whose intensity vanishes proportionally to the volume. It will turn out that the error term, corresponding to the second term in the RHS of (6.7) below, is of smaller order if compared to the remaining terms.

### v3

@@ 5 Probabilistic estimates
The following inequalities and identities hold:
$P_{L,K,\beta,h,r}\big(m_{\{o,z\}}^{i}>M\big)$ $\leq P_{L,K,\beta,h,r}\big(n_{o}-m_{\{o,g\}}^{2}>M\big)$
$\leq\sum\limits_{p=M+1}^{\infty}P_{L,K,\beta,h,r}\big(n_{o}-m_{\{o,g\}}^{2}=p\big)$
$=\sum\limits_{p=M+1}^{\infty}\langle\,\Psi_{p,o}\,\big\rangle$
$\leq\sum\limits_{p=M+1}^{\infty}\Big(\langle\prod_{x\in V}\Psi_{p,x}\big\rangle\Big)^{\frac{1}{|V|}}$
$=\sum\limits_{p=M+1}^{\infty}P_{L,K,\beta,h,r}\big(\forall x\in V:n_{x}-m_{\{x,g\}}^{2}=p\big)^{\frac{1}{|V|}}$
$\leq\sum\limits_{p=M+1}^{\infty}\Big(\frac{1}{Z^{path}_{L,K,\beta,h,r}}d^{|V|}{u(p)}^{|V|/8}\,\,e^{2\,(|E|+|E_{g}|)\max{\{\beta,1\}}}\Big)^{\frac{1}{|V|}}$
$\leq c\sum\limits_{p=M+1}^{\infty}{u(p)}^{\frac{1}{8}}.$
For the first inequality we used that, by definition of local time, Definition 3.5, for $z\in V$ neighbour of $o\in V$, we have that $m^{i}_{o,z}\leq n_{o}-m^{2}_{o,g}$; for the first identity we used Proposition 3.12; for the third inequality we use Proposition 4.9. For the fourth inequality we used a crude upper bound for the probability of the event $\{\forall x\in V:n_{x}-m_{\{x,g\}}^{2}=p\}$. The upper bound is as follows: for each $x$, since $n_{x}-m_{\{x,g\}}^{2}=p$, then, by definition of local time, there exists one of the $d$ edges of $E$ incident to $x$ with at least $\lfloor p/d\rfloor$ dimers of colour $2$ (say, an ‘high occupancy’ edge). Observe that there are at most $d^{V}$ possibilities for choosing a configuration of high occupancy edges and that there are at least $\lfloor\,|V|/8\,\rfloor$ high occupancy edges (at least one for each even vertex). Once one of configuration of high occupancy edges is fixed, we sum over the number of dimers independently for each edge and for each colour, ignoring the constraint that the number of blue dimers equals the number of red dimers at each vertex, using that $U(n)\leq 1$ for each $n\in\mathbb{N}_{0}$, and factorising the sum over the matchings with the factor $\frac{1}{n_{x}(w)!}$ in the definition of the measure, Definition 3.6. We then obtain for each edge in $E$ which is not high occupancy and for each colour a multiplicative factor $e^{\beta}$, for each edge in $E_{g}$ which is not high occupancy a multiplicative factor $e^{h\beta}$, and for each edge which is high occupancy (and is then in $E$) and for the colour $2$ a multiplicative factor at most
$u(p):=\sum\limits_{\ell=\lfloor p/d\rfloor}\frac{\beta^{\ell}}{\ell!},$ (5.6)
thus giving the desired inequality. For the last inequality we used used (5.3). The proof then follows by observing that $\sum_{p=M}^{\infty}{u(p)}^{\frac{1}{8}}$ converges to zero with $M$.
@@
For (5.5) we observe that if $U(0)=0$ then the proof is obvious by definition of the measure, and by the fact that $U(n)>0$ for some $n>0$ by assumption. Hence, we can assume that $U(0)>0$. We have that for some $c\in(0,1)$ we have
$P_{L,K,\beta,h,r}\big(n_{o}=0\big)$ $=\langle\Psi_{o}\big\rangle$
$\leq\Big(\langle\prod_{x\in V}\Psi_{x}\big\rangle\Big)^{\frac{1}{|V}|}$
$=\Big(P_{L,K,\beta,h,r}\big(\forall x\in V,n_{x}=0\big)\Big)^{\frac{1}{|V|}}$
$=\Big(\frac{U(0)^{|V|}}{Z^{path}_{G,\beta,h,r}}\Big)^{\frac{1}{|V|}}\leq c.$
where the first identity follows from Proposition 3.12, the first inequality follows from Proposition 4.9, and the second identity follows again from Proposition 3.12, the third identity follows from the definition of the measure of the multi-occupancy double dimer model, Definition 3.6. In order to see that the last inequality holds for some $c<1$ we need to lower bound $Z^{path}_{G,\beta,h,r}$. For this, let $k>0$ be an integer such that $U(k)>0$. We identify a subset $A\subset E$ of edges such that each pair of edges in $A$ touch distinct vertices and such that $|A|>\frac{1}{100}|V|$. We then lower bound $Z^{path}_{G,\beta,h,r}$ by the weight of the configurations $w\in\mathcal{W}^{\ell}$ such that each edge in $A$ is either empty or it host precisely $k$ blue and $k$ red dimers, while all the remaining edges are empty. This leads to the trivial lower bound
$Z^{path}_{G,\beta,h,r}\geq U(0)^{|V|-2|A|}(U(0)^{2}+\frac{\beta^{2k}U(k)^{2}}{k!^{2}})^{{|A|}}\geq U(0)^{|V|}\big(1+\frac{\beta^{2k}U(k)^{2}}{k!^{2}U(0)^{2}}\big)^{\frac{|V|}{100}}.$
This leads to the desired bound and concludes the proof. ∎
@@
For any $\beta>0$, $h\in[0,1]$, $r\in\mathbb{N}$, there exists a constant $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$, for any $z\in V$ neighbouring $o\in V$ we have that
$\langle S_{o}^{i}S_{z}^{i}\rangle_{L,K,\beta,h,r}\,\,\,$ $=\frac{1}{\beta}E_{L,K,\beta,h,r}\big(\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big)\leq\frac{1}{\beta},$ (5.7)
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle_{L,K,\beta,h,r}$ $\geq c.$ (5.8)
@@
$\langle S_{o}^{i}S_{z}^{i}\rangle_{L,K,\beta,h,r}=\frac{\mu_{\beta,h,r}(\mathcal{A})}{Z^{path}_{L,K,\beta,h,r}},$
where $\mathcal{A}$ is the set of configurations with precisely one $i$-dimer on $\{o,s\}$, one $i$-dimer on $\{z,s\}$, and no further dimer touching $s$. We introduce a map $f$ which takes any configuration $(m,\pi)\in\mathcal{A}$ and associates to it the configuration $f(m,\pi)\in\mathcal{W}^{\ell}$ which is defined by following the next three steps:
@@
    adding a $i$-dimer on $\{o,z\}$,
@@
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle=\sum\limits_{z\sim o}\frac{1}{\beta}E\big(\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big)\geq\sum\limits_{z\sim o}\frac{1}{M\beta}P\big(0<m_{\{o,z\}}^{i},n_{o}\leq M\big)\\ \geq\sum\limits_{z\sim o}\frac{1}{M\beta}\Big(P\big(m_{\{o,z\}}^{i}>0\big)-P\big(n_{o}>M\big)\Big)\\ \geq\frac{1}{M\beta}\Big(P\big(n_{o}>0\big)-6P\big(n_{o}>M\big)\Big)\geq c.$
where the last inequality follows from both inequalities in Lemma 5.2 by choosing $M$ large enough depending on $\beta$. ∎
We now introduce an important quantity of interest.
Definition 5.4 (Magnetisation).
For any $\beta,h,r\geq 0$ we introduce the magnetisation
@@
Set $\beta\geq 0$, $r\in\{1,2\}$, $\ell\in\mathbb{N}$. There exist a constant $c=c(\beta,\ell,r)\in(0,\infty)$ and a value $h_{0}>0$ small enough such that for any $\tilde{h}\in(0,h_{0})$, $L,K\in 2\mathbb{N}\cup\{1\}$
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)$ $\geq\,\,\frac{\beta\tilde{h}}{|\mathbb{S}_{L,K}|}\sum\limits_{x\in\mathbb{S}_{L,K}}\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle_{L,K,\beta,0,r},$ (5.10)
$\langle\cos(\ell s_{o}^{1})\rangle_{L,K,\beta,h_{L,K},r}$ $\leq c\,\tilde{h}$, (5.11)
@@
From a Taylor expansion of the exponentials in (2.5) we deduce that
@@
For the lower bound in (5.10) we use the positive-expectation condition (recall Definition 3.9) and reduce the sum in (5.12) to the terms such that $m_{x}=1$ for just one $x\in V$ and $m_{y}=0$ for all $y\neq x$. We then obtain (5.10) by our choice of $h_{L,K}$ and the condition $\ell=r$. For the upper bound in (5.11) we observe that there exists some $c=c(\beta,\ell,r)\in(1,\infty)$ such that, for each $m$ in the sum (5.12),
$\langle\cos(\ell s_{o}^{1})\prod_{x\in V}{\cos(rs_{x}^{1})}^{m_{x}}\rangle_{L,K,\beta,0,r}\leq c\,\prod_{x\in V}c^{m_{x}}.$
Indeed, the function $\cos(\ell s_{o}^{1})\cos(rs_{o}^{1})^{m_{o}}$ satisfies the assumptions in Lemma 5.1 with $R\leq rm_{o}+\ell$ and $M\leq 2^{m_{o}+1}$ and the function $\cos(rs_{x}^{1})^{m_{x}}$ satisfies the assumptions in Lemma 5.1 with $R\leq rm_{x}$ and $M\,\leq\,2^{m_{x}}$. We then obtain from (5.12) that
@@
where in the last step we used that $\tilde{h}$ is small enough depending on $\beta$, $r$ and $\ell$. ∎
6 Bogoliubov-type inequality
This section presents the central argument for our new proof of the Mermin–Wagner theorem for non positive (but reflection positive) measures.
@@ 6 Bogoliubov-type inequality
$\tilde{H}:=H+\sum\limits_{x\in\mathbb{S}^{e}_{L,k}}\log(\gamma_{x})+\sum\limits_{x\in\mathbb{S}^{o}_{L,k}}\log(\overline{\gamma_{x}}),$
@@
$A:=(\sin(rs_{x}^{1}))_{x\in\mathbb{S}_{L,K}}\qquad\qquad B:=\beta\,\big(\partial_{s_{x}^{1}}\tilde{H}-\partial_{s_{x}^{2}}\tilde{H}\big)_{x\in\mathbb{S}_{L,K}},$
@@
From now on we will sometimes omit the subscripts from $\langle\cdot\rangle$ to simplify the notation. For the first identity we note using integration by parts and the periodicity of the function in the average to conclude that,
@@
[…]
@@ 6.1 Positivity of the Fourier transform
We next state an inequality which is a discrete analogue of the fact that the Fourier transform of a convex, symmetric function on [L,L][-L,L] is nonnegative. The bound will be applied to two-point functions (4.11). Since our convexity property (4.12) fails at a neighbourhood of zero, we obtain a bound which may be negative, but which is negligible comparing to the volume $|\mathbb{S}_{L,K}|$. We define for any $M\in\mathbb{N}$, $\mathbb{Z}^{M}_{odd}:=(2\mathbb{Z}+1)\cap(-M/2,M/2]$ and $\mathbb{Z}^{M}_{even}:=(2\mathbb{Z})\cap(-M/2,M/2]$.
@@
    G(x)=G(x)G(x)=G(-x) for any $x\in e_{1}\mathbb{Z}^{L}_{odd}\cup e_{2}\mathbb{Z}^{L}_{odd}\cup e_{3}\mathbb{Z}^{K}_{odd}$,
@@
$\sum_{x\in e_{i}\mathbb{Z}_{odd}^{L}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{i})-G(5e_{i}))\frac{1-\cos 3k_{i}}{2\sin^{2}(k_{i})},$ (6.3)
$\sum_{x\in e_{3}\mathbb{Z}_{odd}^{K}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{3})-G(5e_{3}))\frac{1-\cos 3k_{3}}{2\sin^{2}(k_{3})},$ (6.4)
@@
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{2\sin^{2}(k_{i})}\sum_{x\in\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x)$ (6.5)
$\sum_{x\in e_{3}\mathbb{Z}^{K}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{2\sin^{2}(k_{3})}\sum_{x\in\mathbb{Z}^{K}_{odd}}(2G(x)-G(x-2e_{3})-G(x+2e_{3}))\cos(k\cdot x).$ (6.6)
@@
$=\frac{1}{2\sin^{2}(k_{i})}\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x),$
@@
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $\geq\frac{1}{2\sin^{2}(k_{1})^{2}}\big(2(G(5e_{i})-G(3e_{i}))+2(2G(3e_{i})-G(e_{i})-G(5e_{i}))\cos(3k_{i})$
@@
$=\frac{1}{2\sin^{2}(k_{1})}\left((G(e_{i})-G(3e_{i}))(\cos(k_{i})-\cos(3k_{i}))+(G(3e_{i})-G(5e_{i}))(\cos(3k_{i})-1)\right),$
@@
6.2 Statement and proof of the Bogoliubov-type inequality
We now state a Bogoliubov-type inequality for the quantities $A$ and $B$ which have been defined above. Our inequality holds for small external magnetic fields, namely whose intensity vanishes proportionally to the volume. It will turn out that the error term, corresponding to the second term in the RHS of (6.7) below, is of smaller order if compared to the remaining terms.